# Higher-rank GBS groups: non-positive curvature and biautomaticity

### v1

@@ 1. Introduction
In this note, we generalise the results of Leary and Minasyan to all rank $n$ generalised Baumslag–Solitar groups (or rank $n$ GBS groups for short)—fundamental groups of (finite) graphs of groups ${\mathcal{G}}$ with all vertex and edge groups isomorphic to ${\mathbb{Z}}^{n}$ for a fixed $n\in{\mathbb{N}}$. (We call such a graph of groups ${\mathcal{G}}$ a rank $n$ GBS graph of groups.) Our criteria are based on the image of the modular homomorphism $\Delta_{\mathcal{G}}$ as a subgroup of the abstract commensurator $\operatorname{Comm}(H)\cong\operatorname{GL}_{n}({\mathbb{Q}})$, where $H$ is a vertex group of ${\mathcal{G}}$. See Section 2 for the precise terminology and notation.

### v2

@@ 1. Introduction
In this note, we generalise the results of Leary and Minasyan to all rank $n$ generalised Baumslag–Solitar groups (or rank $n$ GBS groups for short)—fundamental groups of (finite) graphs of groups ${\mathcal{G}}$ with all vertex and edge groups isomorphic to ${\mathbb{Z}}^{n}$ for a fixed $n\in{\mathbb{N}}$. (We call such a graph of groups ${\mathcal{G}}$ a rank $n$ GBS graph of groups.) Rank 1 GBS groups are well studied, for instance from the point of view of splittings, quasi-isometries, automorphisms and commensurability [Why01, For03, For06, Lev07, Lev15, CRKZ21]. The general setting of rank $n$ GBS groups has also been studied by several authors in recent years, from the perspective of actions on hyperbolic spaces in [But22] and the perspective of separability properties in [LdGZS23].
The criteria for our theorem are based on the image of the modular homomorphism $\Delta_{\mathcal{G}}$ as a subgroup of the abstract commensurator $\operatorname{Comm}(H)\cong\operatorname{GL}_{n}({\mathbb{Q}})$, where $H$ is a vertex group of ${\mathcal{G}}$. See Section 2 for the precise terminology and notation.
@@
After writing the first version of this paper, we learnt that part (ii) of Theorem 1.1 had been independently proved by Button and was due to appear in his upcoming publication [But22].